# Efficient GW band structure calculations using Gaussian basis functions and application to atomically thin transition-metal dichalcogenides

### v2

@@ V.3 GW space-time calculations (CP2K)
We have implemented the low-scaling $GW$ space-time algorithm presented in this work in the CP2K software package [82, 83, 84]. CP2K employs a Gaussian basis set for representing KS orbitals [Eq. (17)] and a plane-waves basis set for the electron density to evaluate the Hartree potential via Ewald summation [82, 83]. We use Gaussian dual-space pseudopotentials [54]. In the DFT calculation, we employ the PBE exchange-correlation functional [85]. The plane-wave cutoff for the electron density is set to $500$ Ry. This value was converged beforehand on the DFT-level results, as the plane-wave grid is not used for the $GW$ part of the calculations.
@@
In the $GW$ algorithm, we employ a minimax time-frequency grid [49, 70, 71]. We compute two- and three-centre integrals over Gaussians using analytical schemes [86, 87]. The self-energy is analytically continued from imaginary frequency to the real frequency using a Padé model [88, 6, 60] with 16 parameters. Unless otherwise noted, we employ a cutoff radius $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt7$ Å, an RI regularization $\alpha\hskip 1.69998pt{=}\hskip 1.69998pt10^{-2}$ and a box height for the 2D materials of 15 Å (for computing the Fourier transform in the Hartree potential). As already stated, the extrapolation of the $k$-integration of $W$ [Eq. (51)] relies on two $k$-meshes, one 4 times denser than the equivalent DFT $k$-mesh along each direction and one 8 times denser (e.g. for a $32\hskip 0.89996pt{\times}\hskip 0.59998pt32$ DFT $k$-mesh, the corresponding coarse W $k$-mesh is $128\hskip 0.89996pt{\times}\hskip 0.59998pt128$ and the dense $k$-mesh is $256\hskip 0.89996pt{\times}\hskip 0.59998pt256$). The remaining computational parameters include the number of minimax time and frequency points $N_{\tau/\omega}$ [49, 70, 71], the DFT $k$-mesh [Eqs. (21), (23), (32); the DFT $k$-mesh also defines the SC supercell (29)], and the filter threshold for sparse matrix-tensor operations in Eqs. (42), (56) and (59). In this paper, we will use two sets of parameters: tight settings ($N_{\tau/\omega}\hskip 1.69998pt{=}\hskip 1.69998pt30$, DFT $k$-mesh: $32\hskip 0.89996pt{\times}\hskip 0.59998pt32$, filter: $10^{-11}$) which correspond to a reference set of parameters that we have defined through extensive testing in order to yield well-converged $GW$ band gaps, and a set of light settings ($N_{\tau/\omega}\hskip 1.69998pt{=}\hskip 1.69998pt10$, DFT $k$-mesh: $24\hskip 0.89996pt{\times}\hskip 0.59998pt24$, filter: $10^{-6}$) that have sufficiently reduced memory and computational costs in order to be used for laptop calculations while still giving decently accurate results (quantitative values are given in Sec. VI and Sec. VIII). Note that the memory bottleneck comes from the parallelization strategy, as one can see in Appendix C, which implies the storage of all the three-centre integrals on each parallel group. In theory, one could also recompute these at each loop of the code, which would in practice prohibitively increase the computational time.
@@
We use the single-, double- and triple-zeta MOLOPT basis sets [89] for expanding the KS orbitals, Eq. (17). These basis sets have been optimized for the total energy of the ground state so that they might exhibit a slow convergence behaviour for excited states, and therefore for band gaps at the $GW$ level. This motivated us to also use augmented single-, double- and triple-zeta Gaussian basis sets (aug-SZV-MOLOPT, aug-DZVP-MOLOPT and aug-TZVP-MOLOPT). We have created those by augmenting Gaussian SZV-MOLOPT, DZVP-MOLOPT and TZVP-MOLOPT bases [89] of S, Se, Mo, W with an additional $s$, $p$, $d$ function, an additional $f$ function (for Mo, W and aug-DZVP-MOLOPT and aug-TZVP-MOLOPT of S, Se) and an additional $g$ function (all aug-TZVP-MOLOPT and aug-DZVP-MOLOPT of Mo, W). We have optimized the contraction coefficients of the additional functions by optimizing the lowest five GWGW+Bethe-Salpeter excitation energies [90] of a molecular set [89]. We will report DFT and $GW$ with the original, i.e., non-augmented MOLOPT basis sets and the augmented MOLOPT basis sets.
@@
For the density response (42), dielectric function (44) and screened Coulomb interaction (47), an auxiliary Gaussian RI basis set is required. There is no general approach for constructing optimally sized RI basis sets, as their design is closely tied to the chosen AO basis. In the case of the widely used cc-pVNZ Dunning basis sets [91], for example, the corresponding RI basis sets introduced in Ref. [92] provide a single, fixed RI basis for each AO basis. As a result, convergence with respect to the size of the RI basis is not commonly investigated. It is also possible to generate RI basis sets on the fly [93], although this usually leads to fairly large numbers of basis functions. In our case, we have optimized the RI basis set $\{\varphi_{P}\}$ by matching the RI-MP2 correlation energy [94] of single atoms to the MP2 correlation energy. Unless stated otherwise, all reference calculations in this paper were carried using an RI basis with a relative RI-MP2 correlation energy difference of $10^{-3}$ with respect to the corresponding MP2 correlation energy.
@@
We incorporate SOC via parameters from HGH Gaussian dual-space pseudopotentials [54, 55, 56], see Eqs. (31), (64) and Appendix H for details. We compute the SOC for states in a window of 40 eV, so 20 eV below the valence band maximum to 20 eV above the conduction band minimum, in order to avoid possible numerical instabilities with bands far from the gap. For the case of WSe2, we chose a window to 20 eV for aug-DZVP-MOLOPT and aug-TZVP-MOLOPT, as discussed in Appendix H.
@@ V.4 Reference GW calculations with a plane-wave-based algorithm (BerkeleyGW)
We performed the Quantum Espresso (QE) [95] DFT calculations employing the PBE exchange-correlation functional [85]. Fully relativistic norm-conserving pseudopotentials were used, as provided by the PseudoDojo database [96]. A plane-wave energy cut-off of 100 Ry was applied, and the self-consistent charge density was converged on a $30\times 30\times 1$ $\mathbf{k}$-grid with a total energy convergence threshold of $10^{-9}$ Ry.
@@
Using the QE DFT energies and states, we performed for each material, a one-shot $GW$ calculation ($G_{0}W_{0}$) using the BerkeleyGW package [97, 57]. We considered the full spinor implementation of BerkeleyGW [98], which incorporates SOC non-perturbatively. The dielectric matrix was computed with a dielectric cut-off of 25 Ry, considering a total of 3999 occupied and empty bands on a $12\times 12\times 1$ uniform $\mathbf{k}$-grid. For completeness, we computed the quasi-particle band-gap using the generalized plasmon-pole model of Hybertsen-Louie [97] (see Appendix I) and the full-frequency evaluation of the self-energy. In the explicit full frequency calculation, we used the contour-deformation method with the Adler-Wiser formula. We employed a frequency step and broadening of 0.25 eV, using 15 frequency points along the imaginary axis within the contour deformation approach. A low-frequency cutoff of 20.0 eV was set to restrict the real-axis integration range. To accelerate convergence in the vicinity of $|\mathbf{q}|\rightarrow 0$, a non-uniform neck subsampling approach was employed [99] and the spurious interactions between periodic replicas in the perpendicular direction to the surface were removed with a Coulomb interaction truncation scheme [100]. The full frequency dependence of self-energy was evaluated with a frequency step of $0.2$ eV in the frequency range [2.0,2.0][-2.0,2.0] eV and centering the frequency grid around each mean-field quasi-particle energy. For the band structures, the quasi-particle energies computed on the coarse $\mathbf{k}$-grid were interpolated along high-symmetry $\mathbf{k}$-paths.
@@ V.5 Reference GW calculations with the projector-augmented-wave scheme (VASP)
To obtain the quasi-particle band gap, we carry out a single-shot $G_{0}W_{0}$ calculation using the Vienna Ab initio Simulation Package (VASP) [101, 102]. In this process, the initial KS wavefunctions and energy levels, derived from a preceding DFT calculation also performed with VASP, are used. Core electrons are treated using the $GW$ adaptation of the projector-augmented-wave (PAW) pseudopotentials [103, 104]. For all TMDCs, we apply PAWs constructed on the Perdew–Burke–Ernzerhof (PBE) functional [85] with an energy cutoff of 500 eV. The considered valence electron configurations are $4s^{2}4p^{6}5s^{1}4d^{5}$ for Mo, $3s^{2}3p^{4}$ for S, $5p^{6}6s^{2}5d^{4}$ for W, $4s^{2}4p^{4}$ for Se. A total of 384 bands, comprising both occupied and unoccupied states, are taken into account, along with a uniform $\mathbf{k}$-grid of $18\times 18\times 1$ for MoS2 and WS2, and of $15\times 15\times 1$ for MoSe2 and WSe2, ensuring smooth convergence in the vicinity of $|\mathbf{q}|\rightarrow 0$. Moreover, the dielectric tensor is broadened with a Lorentzian of 0.01 eV in all cases except for MoSe2, where a broadening of 0.1 eV is applied. Finally, the self-energy is calculated using a full-frequency implementation with 100 points along the imaginary frequency grid, as provided in VASP.
@@
We show the DFT band structure computed with the PBE exchange-correlation functional [85] and SOC in Fig. 1, for the aug-SZV-MOLOPT basis set (light settings) and the aug-DZVP-MOLOPT (tight settings), and the band structure computed from QE (with 8 empty bands). We observe excellent agreement between the aug-DZVP-MOLOPT calculation and the QE calculation, with a difference of 18 meV on average between their respective PBE+SOC bandgaps (see Table 1), and also with VASP results with a difference of 10 meV. The agreement of the PBE+SOC gap computed with the small aug-SZV-MOLOPT basis is also good, the average deviation is 29 meV to QE and 23 meV to VASP (Table 1). This shows a good agreement of DFT band structures between each code, which validates their use as a starting point for a $GW$ benchmark. Table 1 also reports the DFT band gaps without SOC and the SOC splitting at the K-point, demonstrating that our approach yields SOC splittings in good agreement with plane-wave reference calculations.
@@
For comparison, we also computed the band gaps using the original, non-augmented MOLOPT basis sets [89] (see Table 1). At the DFT level, the results show good agreement with those from augmented basis sets—except for SZV-MOLOPT—indicating that augmentation is not strictly required for fast convergence of the DFT band gap with respect to basis set size.
@@
To assess the accuracy of the SOC implementation, we report in Table 1 the spin-orbit splitting of the valence band maximum at the K-point. This splitting is crucial as it is involved in determining the (optical) energy difference between A and B excitons, observed in reflectance and photoluminescence spectra [105, 106]. It thus plays a crucial role for valley selective optical excitations. We observe good agreement between our perturbative SOC implementation and the fully relativistic implementation in QE, with an average difference of 17 meV between the aug-DZVP-MOLOPT and QE results, as an example. The difference is more significant for the selenium-based TMDs, with 32 meV on average whereas the difference is 2 meV for the sulfur-based TMDs. The results are similar for the comparison with the VASP SOC-splitting. This finding validates our perturbative SOC treatment from HGH pseudopotentials [Eqs. (31), (64) and Appendix H].
@@
We demonstrate the impact of SOC on the band structure by giving in Appendix H the comparison between the PBE and PBE+SOC bandstructures for all monolayers (Fig. 10). One can therefore see that the SOC lifts the spin degeneracy and therefore splits the band structure, especially at the K-point as we have already discussed, This leads to a reduction of the band gap with respect to the calculation without SOC, especially in the case of WS2 and WSe2 where the band gap is lowered by 0.3 eV (Table 1).
@@
Figure 2: $G_{0}W_{0}$@PBE+SOC band gap of monolayer WSe2 and execution time as a function of the number of time points $\tau$ (Sec. IV), the $k$-mesh (Eqs. (21), (23), (32)), the filter threshold (for Eqs. (42), (56) and (59), see Eqs. (88) and (89)), the simulation cell box height (Sec. V.3), the size factor $\Delta$ for $V_{PQ}(\mathbf{k})$ (Eq. (110)) and the number of basis functions (Sec. V.3). Default parameters are reported on top. In (a), we also show $G_{0}W_{0}$@PBE+SOC with Hedin’s shift [1, 107, 108, 109, 3, 110] to avoid poles of the self-energy close to the quasiparticle solution [111, 112].
@@
Figure 3: Direct $G_{0}W_{0}$@PBE+SOC band gap of monolayer WSe2 as function of the RI basis set size, the truncation radius $r_{\text{c}}$ of the truncated Coulomb metric (37)/(41) and the regularization parameter $\alpha$, Eq. (45). Tight parameters and the aug-SZV-MOLOPT basis sets are used to expand the KS orbitals. Each point corresponds to an RI basis set with a given relative deviation of the RI-MP2 [94] correlation energy compared to MP2: the RI basis set with $N_{\text{RI}}\hskip 1.69998pt{=}\hskip 1.69998pt88$ functions gives a relative RI-MP2 error of $10^{-2}$; the other RI basis set sizes and relative RI-MP2 errors are: $N_{\text{RI}}=130$ and $10^{-3}$, $N_{\text{RI}}=141$ and $10^{-4}$, $N_{\text{RI}}=187$ and $10^{-5}$.
@@
Using our tight convergence settings (Sec. V.3), we obtain a $G_{0}W_{0}$@PBE+SOC band gap of 1.95 eV with the aug-SZV-MOLOPT basis and 1.93 eV with the aug-DZVP-MOLOPT basis (Table 1). A larger augmented basis set is not used due to current memory limitations in our implementation. Fig. 2 shows how the band gap changes with various computational parameters. In Fig. 2a, we test different sizes of the time-frequency grid. We find that the gap changes by up to 12 meV in $G_{0}W_{0}$@PBE (blue trace), when using 14, 20, 26, 30 and 34 points. These variations can arise by poles in the self-energy [112], which is an unphysical feature of $G_{0}W_{0}$@PBE and eigenvalue self-consistency in $G$ will cure this deficiency [111, 112]. We apply Hedin’s shift [1, 107, 108, 109, 3, 110] to approximate eigenvalue self-consistency in $G$, which reduces the variation between 14 and 34 time-frequency points to 3 meV (red curve in Fig. 1a).
@@
Note that the number of cells included in the lattice sums (42), (56) and (59) depend on both the filter threshold of three-centre integrals (Fig. 2e) and the supercell "SC" defined in Eq. (29). The size of the supercell "SC" is determined by the DFT $k$-mesh [Eq. (29)]. We thus demonstrate convergence w.r.t. the number of cells in the lattice sums (42), (56) and (59) by demonstrating convergence with the $k$-mesh (Fig. 2c) and the filter threshold (Fig. 2e). Similarly, we show the fast convergence of the lattice sum (46) of the Coulomb matrix in Fig. 2i (the size factor of the lattice summation is defined in Appendix D). Also, the box height (Fig. 2g) can be well-converged.
@@
We also study the convergence of the $GW$ gap with respect to the RI basis size in Fig. 3. Without regularization $(\alpha\hskip 1.69998pt{=}\hskip 1.69998pt0$), the band gap increases with RI basis size (Fig. 3a). We assign this issue to linear dependencies in the RI basis set, which leads to numerical instabilities in the RI basis expansion, see Appendix A, Eq. (66). In a nutshell, numerical instabilities arise when two spatially close diffuse $s$-type RI function can partially compensate each other, leading to large expansion coefficients $B_{P\mathbf{P}}^{\mu\mathbf{R}\nu\mathbf{T}}$ in Eq. (66) with alternating sign.
To fix this issue, we use RI regularization [47] via $\alpha\hskip 1.69998pt{=}\hskip 1.69998pt10^{-3}$ and $\alpha\hskip 1.69998pt{=}\hskip 1.69998pt10^{-2}$ in Eq. (45). RI regularization ensures numerical stability, see Fig. 3b,c: For an RI basis set size of a single WSe2 unit between 130 and 187 and a cutoff radius $r_{\text{c}}\hskip 1.69998pt{\in}\hskip 1.69998pt\{5\,\text{\AA },7\,\text{\AA },9$ Å}, the $GW$ gap is identical within 27 meV.

### v3

@@ V.3 GW space-time calculations (CP2K)
We have implemented the low-scaling $GW$ space-time algorithm presented in this work in the CP2K software package [Kuehne2020, Kuehne2025, cp2k]. CP2K employs a Gaussian basis set for representing KS orbitals [Eq. (17)] and a plane-wave basis set for the electron density to evaluate the Hartree potential via Ewald summation [Kuehne2020, Kuehne2025]. We use Gaussian dual-space pseudopotentials [Hartwigsen1998]. In the DFT calculation, we employ the PBE exchange-correlation functional [Perdew1996]. The plane-wave cutoff for the electron density is set to $500$ Ry. This value was converged beforehand on the DFT-level results, as the plane-wave grid is not used for the $GW$ part of the calculations.
@@
In the $GW$ algorithm, we employ a minimax time-frequency grid [liu2016, Azizi2023, Azizi2024]. We compute two- and three-centre integrals over Gaussians using analytical schemes [Golze2017, libint]. The self-energy is analytically continued from imaginary frequency to the real frequency using a Padé model [Vidberg1977, gw100, Leucke2025] with 16 parameters. Unless otherwise noted, we employ a cutoff radius $r_{c}\hskip 1.69998pt{=}\hskip 1.69998pt7$ Å, an RI regularization $\alpha\hskip 1.69998pt{=}\hskip 1.69998pt10^{-2}$ and a box height for the 2D materials of 15 Å (for computing the Fourier transform in the Hartree potential). As already stated, the extrapolation of the $k$-integration of $W$ [Eq. (51)] relies on two $k$-meshes, one 4 times denser than the equivalent DFT $k$-mesh along each direction and one 8 times denser (e.g. for a $32\hskip 0.89996pt{\times}\hskip 0.59998pt32$ DFT $k$-mesh, the corresponding coarse W $k$-mesh is $128\hskip 0.89996pt{\times}\hskip 0.59998pt128$ and the dense $k$-mesh is $256\hskip 0.89996pt{\times}\hskip 0.59998pt256$). The remaining computational parameters include the number of minimax time and frequency points $N_{\tau/\omega}$ [liu2016, Azizi2023, Azizi2024], the DFT $k$-mesh [Eqs. (21), (23), (32); the DFT $k$-mesh also defines the SC supercell (29)], and the filter threshold for sparse matrix-tensor operations in Eqs. (42), (56) and (59). In this paper, we will use two sets of parameters: tight settings ($N_{\tau/\omega}\hskip 1.69998pt{=}\hskip 1.69998pt30$, DFT $k$-mesh: $32\hskip 0.89996pt{\times}\hskip 0.59998pt32$, filter: $10^{-11}$) which correspond to a reference set of parameters that we have defined through extensive testing in order to yield well-converged $GW$ band gaps, and a set of light settings ($N_{\tau/\omega}\hskip 1.69998pt{=}\hskip 1.69998pt10$, DFT $k$-mesh: $24\hskip 0.89996pt{\times}\hskip 0.59998pt24$, filter: $10^{-6}$) that have sufficiently reduced memory and computational costs in order to be used for laptop calculations while still giving decently accurate results (quantitative values are given in Sec. VI and Sec. VIII). Note that the memory bottleneck comes from the parallelization strategy, as one can see in Appendix C, which implies the storage of all the three-centre integrals on each parallel group. In theory, one could also recompute these at each loop of the code, which would in practice prohibitively increase the computation time.
@@
We use the single-, double- and triple-zeta MOLOPT basis sets [vande2007] for expanding the KS orbitals, Eq. (17). These basis sets have been optimized for the total energy of the ground state so that they might exhibit a slow convergence behaviour for excited states, and therefore for band gaps at the $GW$ level. This motivated us to also use augmented single-, double- and triple-zeta Gaussian basis sets (aug-SZV-MOLOPT, aug-DZVP-MOLOPT and aug-TZVP-MOLOPT). We have created those by augmenting Gaussian SZV-MOLOPT, DZVP-MOLOPT and TZVP-MOLOPT bases [vande2007] of S, Se, Mo, W with an additional $s$, $p$, $d$ function, an additional $f$ function (for Mo, W and aug-DZVP-MOLOPT and aug-TZVP-MOLOPT of S, Se) and an additional $g$ function (all aug-TZVP-MOLOPT and aug-DZVP-MOLOPT of Mo, W). We have optimized the contraction coefficients of the additional functions by optimizing the lowest five GWGW+Bethe-Salpeter excitation energies [Blase2020] of a molecular set [vande2007]. We will report DFT and $GW$ with the original, i.e., non-augmented MOLOPT basis sets and the augmented MOLOPT basis sets.
@@
For the density response (42), dielectric function (44) and screened Coulomb interaction (47), an auxiliary Gaussian RI basis set is required. There is no general approach for constructing optimally sized RI basis sets, as their design is closely tied to the chosen AO basis. In the case of the widely used cc-pVNZ Dunning basis sets [Dunning1989], for example, the corresponding RI basis sets introduced in Ref. [Weigend2002] provide a single, fixed RI basis for each AO basis. As a result, convergence with respect to the size of the RI basis is not commonly investigated. It is also possible to generate RI basis sets on the fly [Stoychev2017], although this usually leads to fairly large numbers of basis functions. In our case, we have optimized the RI basis set $\{\varphi_{P}\}$ by matching the RI-MP2 correlation energy [Weigend1998] of single atoms to the MP2 correlation energy. Unless stated otherwise, all reference calculations in this paper were carried out using an RI basis with a relative RI-MP2 correlation energy difference of $10^{-3}$ with respect to the corresponding MP2 correlation energy.
@@
We incorporate SOC via parameters from HGH Gaussian dual-space pseudopotentials [Hartwigsen1998, Krack2005, Vogt2025], see Eqs. (31), (64) and Appendix H for details. We compute the SOC for states in a window of 40 eV, so 20 eV below the valence band maximum to 20 eV above the conduction band minimum, in order to avoid possible numerical instabilities with bands far from the gap. For the case of WSe2, we chose a window of 20 eV for aug-DZVP-MOLOPT and aug-TZVP-MOLOPT, as discussed in Appendix H.
@@ V.4 Reference GW calculations with a plane-wave-based algorithm (BerkeleyGW)
We performed the Quantum Espresso (QE) [Giannozzi2017] DFT calculations employing the PBE exchange-correlation functional [Perdew1996]. Fully relativistic norm-conserving pseudopotentials were used, as provided by the PseudoDojo database [Dojo2019]. A plane-wave energy cut-off of 100 Ry was applied, and the self-consistent charge density was converged on a $30\times 30\times 1$ $\mathbf{k}$-grid with a total energy convergence threshold of $10^{-9}$ Ry.
@@
Using the QE DFT energies and states, we performed for each material a one-shot $GW$ calculation ($G_{0}W_{0}$) using the BerkeleyGW package [Hybertsen1986, Deslippe2012]. We considered the full spinor implementation of BerkeleyGW [Barker2022], which incorporates SOC non-perturbatively. The dielectric matrix was computed with a dielectric cut-off of 25 Ry, considering a total of 3999 occupied and empty bands on a $12\times 12\times 1$ uniform $\mathbf{k}$-grid. For completeness, we computed the quasi-particle band-gap using the generalized plasmon-pole model of Hybertsen-Louie [Hybertsen1986] (see Appendix I) and the full-frequency evaluation of the self-energy. In the explicit full frequency calculation, we used the contour-deformation method with the Adler-Wiser formula. We employed a frequency step and broadening of 0.25 eV, using 15 frequency points along the imaginary axis within the contour deformation approach. A low-frequency cutoff of 20.0 eV was set to restrict the real-axis integration range. To accelerate convergence in the vicinity of $|\mathbf{q}|\rightarrow 0$, a non-uniform neck subsampling approach was employed [Qiu2017] and the spurious interactions between periodic replicas in the perpendicular direction to the surface were removed with a Coulomb interaction truncation scheme [Ismail2006]. The full frequency dependence of the self-energy was evaluated with a frequency step of $0.2$ eV in the frequency range [2.0,2.0][-2.0,2.0] eV and centering the frequency grid around each mean-field quasi-particle energy. For the band structures, the quasi-particle energies computed on the coarse $\mathbf{k}$-grid were interpolated along high-symmetry $\mathbf{k}$-paths.
@@ V.5 Reference GW calculations with the projector-augmented-wave scheme (VASP)
To obtain the quasi-particle band gap, we carry out a single-shot $G_{0}W_{0}$ calculation using the Vienna Ab initio Simulation Package (VASP) [vasp1, vasp2]. In this process, the initial KS wavefunctions and energy levels, derived from a preceding DFT calculation also performed with VASP, are used. Core electrons are treated using the $GW$ adaptation of the projector-augmented-wave (PAW) pseudopotentials [PAW1, PAW2]. For all TMDCs, we apply PAWs constructed on the Perdew–Burke–Ernzerhof (PBE) functional [Perdew1996] with an energy cutoff of 500 eV. The considered valence electron configurations are $4s^{2}4p^{6}5s^{1}4d^{5}$ for Mo, $3s^{2}3p^{4}$ for S, $5p^{6}6s^{2}5d^{4}$ for W, $4s^{2}4p^{4}$ for Se. A total of 384 bands, comprising both occupied and unoccupied states, are taken into account, along with a uniform $\mathbf{k}$-grid of $18\times 18\times 1$ for MoS2 and WS2, and of $15\times 15\times 1$ for MoSe2 and WSe2, ensuring smooth convergence in the vicinity of $|\mathbf{q}|\rightarrow 0$. Moreover, the dielectric tensor is broadened with a Lorentzian of 0.01 eV in all cases except for MoSe2, where a broadening of 0.1 eV is applied. Finally, the self-energy is calculated using a full-frequency implementation with 100 points along the imaginary frequency grid, as provided in VASP.
@@
We show the DFT band structure computed with the PBE exchange-correlation functional [Perdew1996] and SOC in Fig. 1, for the aug-SZV-MOLOPT basis set (light settings) and the aug-DZVP-MOLOPT (tight settings), and the band structure computed from QE (with 8 empty bands). We observe excellent agreement between the aug-DZVP-MOLOPT calculation and the QE calculation, with a difference of 18 meV on average between their respective PBE+SOC bandgaps (see Table 1), and also with VASP results with a difference of 10 meV. The agreement of the PBE+SOC gap computed with the small aug-SZV-MOLOPT basis is also good, the average deviation is 29 meV to QE and 23 meV to VASP (Table 1). This shows a good agreement of DFT band structures between each code, which validates their use as a starting point for a $GW$ benchmark. Table 1 also reports the DFT band gaps without SOC and the SOC splitting at the K-point, demonstrating that our approach yields SOC splittings in good agreement with plane-wave reference calculations.
@@
For comparison, we also computed the band gaps using the original, non-augmented MOLOPT basis sets [vande2007] (see Table 1). At the DFT level, the results show good agreement with those from augmented basis sets—except for SZV-MOLOPT—indicating that augmentation is not strictly required for fast convergence of the DFT band gap with respect to basis set size.
@@
To assess the accuracy of the SOC implementation, we report in Table 1 the spin-orbit splitting of the valence band maximum at the K-point. This splitting is crucial as it is involved in determining the (optical) energy difference between A and B excitons, observed in reflectance and photoluminescence spectra [Heinz2010, Urbaszek2018]. It thus plays a crucial role for valley selective optical excitations. We observe good agreement between our perturbative SOC implementation and the fully relativistic implementation in QE, with an average difference of 17 meV between the aug-DZVP-MOLOPT and QE results, as an example. The difference is more significant for the selenium-based TMDs, with 32 meV on average whereas the difference is 2 meV for the sulfur-based TMDs. The results are similar for the comparison with the VASP SOC-splitting. This finding validates our perturbative SOC treatment from HGH pseudopotentials [Eqs. (31), (64) and Appendix H].
@@
We demonstrate the impact of SOC on the band structure by giving in Appendix H the comparison between the PBE and PBE+SOC bandstructures for all monolayers (Fig. 10). One can therefore see that the SOC lifts the spin degeneracy and therefore splits the band structure, especially at the K-point as we have already discussed. This leads to a reduction of the band gap with respect to the calculation without SOC, especially in the case of WS2 and WSe2 where the band gap is lowered by 0.3 eV (Table 1).
@@
Figure 2: $G_{0}W_{0}$@PBE+SOC band gap of monolayer WSe2 and execution time as a function of the number of time points $\tau$ (Sec. IV), the $k$-mesh (Eqs. (21), (23), (32)), the filter threshold (for Eqs. (42), (56) and (59), see Eqs. (88) and (89)), the simulation cell box height (Sec. V.3), the size factor $\Delta$ for $V_{PQ}(\mathbf{k})$ (Eq. (110)) and the number of basis functions (Sec. V.3). Default parameters are reported on top. In (a), we also show $G_{0}W_{0}$@PBE+SOC with Hedin’s shift [hedin1965new, Hedin1999, Pollehn1998, Martin_Reining_Ceperley_2016, Golze2019, Golze2022] to avoid poles of the self-energy close to the quasiparticle solution [Veril2018, Schambeck2024].
@@
Figure 3: Direct $G_{0}W_{0}$@PBE+SOC band gap of monolayer WSe2 as function of the RI basis set size, the truncation radius $r_{\text{c}}$ of the truncated Coulomb metric (37)/(41) and the regularization parameter $\alpha$, Eq. (45). Tight parameters and the aug-SZV-MOLOPT basis sets are used to expand the KS orbitals. Each point corresponds to an RI basis set with a given relative deviation of the RI-MP2 [Weigend1998] correlation energy compared to MP2: the RI basis set with $N_{\text{RI}}\hskip 1.69998pt{=}\hskip 1.69998pt88$ functions gives a relative RI-MP2 error of $10^{-2}$; the other RI basis set sizes and relative RI-MP2 errors are: $N_{\text{RI}}=130$ and $10^{-3}$, $N_{\text{RI}}=141$ and $10^{-4}$, $N_{\text{RI}}=187$ and $10^{-5}$.
@@
Using our tight convergence settings (Sec. V.3), we obtain a $G_{0}W_{0}$@PBE+SOC band gap of 1.95 eV with the aug-SZV-MOLOPT basis and 1.93 eV with the aug-DZVP-MOLOPT basis (Table 1). A larger augmented basis set is not used due to current memory limitations of our implementation. Fig. 2 shows how the band gap changes with various computational parameters. In Fig. 2a, we test different sizes of the time-frequency grid. We find that the gap changes by up to 12 meV in $G_{0}W_{0}$@PBE (blue trace), when using 14, 20, 26, 30 and 34 points. These variations can arise by poles in the self-energy [Schambeck2024], which is an unphysical feature of $G_{0}W_{0}$@PBE and eigenvalue self-consistency in $G$ will cure this deficiency [Veril2018, Schambeck2024]. We apply Hedin’s shift [hedin1965new, Hedin1999, Pollehn1998, Martin_Reining_Ceperley_2016, Golze2019, Golze2022] to approximate eigenvalue self-consistency in $G$, which reduces the variation between 14 and 34 time-frequency points to 3 meV (red curve in Fig. 1a).
@@
Note that the number of cells included in the lattice sums (42), (56) and (59) depend on both the filter threshold of three-centre integrals (Fig. 2e) and the supercell "SC" defined in Eq. (29). The size of the supercell "SC" is determined by the DFT $k$-mesh [Eq. (29)]. We thus demonstrate convergence w.r.t. the number of cells in the lattice sums (42), (56), and (59) by demonstrating convergence with the $k$-mesh (Fig. 2c) and the filter threshold (Fig. 2e). Similarly, we show the fast convergence of the lattice sum (46) of the Coulomb matrix in Fig. 2i (the size factor of the lattice summation is defined in Appendix D). Also, the box height (Fig. 2g) can be well-converged.
@@
We also study the convergence of the $GW$ gap with respect to the RI basis size in Fig. 3. Without regularization $(\alpha\hskip 1.69998pt{=}\hskip 1.69998pt0$), the band gap increases with RI basis size (Fig. 3a). We assign this issue to linear dependencies in the RI basis set, which lead to numerical instabilities in the RI basis expansion, see Appendix A, Eq. (66). In a nutshell, numerical instabilities arise when two spatially close diffuse $s$-type RI functions partially compensate each other, leading to large expansion coefficients $B_{P\mathbf{P}}^{\mu\mathbf{R}\nu\mathbf{T}}$ in Eq. (66) with alternating sign.
To fix this issue, we use RI regularization [Graml2024] via $\alpha\hskip 1.69998pt{=}\hskip 1.69998pt10^{-3}$ and $\alpha\hskip 1.69998pt{=}\hskip 1.69998pt10^{-2}$ in Eq. (45). RI regularization ensures numerical stability, see Fig. 3b,c: For an RI basis set size of a single WSe2 unit between 130 and 187 and a cutoff radius $r_{\text{c}}\hskip 1.69998pt{\in}\hskip 1.69998pt\{5\,\text{\AA },7\,\text{\AA },9$ Å}, the $GW$ gap is identical within 27 meV.